OpenStudy (anonymous):

Marie, Brad, and Doug went to the mall. Brad spent three times as much money as Marie. Doug spent fifteen dollars more than Brad. If Doug spent $60 at the mall, which of the following equations should be used to find out how much money Marie spent? 3(m + 15) = 60 (where m represents the amount of money Marie spent) 3m + 15 = 60 (where m represents the amount of money Marie spent) 15m + 3 = 60 (where m represents the amount of money Marie spent) 3m – 15 = 60 (where m represents the amount of money Marie spent)

OpenStudy (anonymous):

B = 3M D = B + 15 If D spent 60, how much did M spent? D = B + 15 Replace that B with the B from the first equation D = 3M + 15 Since D spent 60, 60 = 3M + 15 Which is option B